# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 5.3 Reconstructing Waddington Developmental Landscapes
Waddington’s landscape metaphor is a widely recognized framework to depict the cell fate decision process. This conceptual model suggests that metastable cellular states are analogous to wells within a potential landscape, and transitions between these states can be understood as movements or “hops” between these potential wells. While the development of such potential landscapes has been extensively explored (Zhou and Li, 2016; Shi et al., 2022; Zhao et al., 2024; Li and Wang, 2013; Wang et al., 2010; Li and Wang, 2014; Bian et al., 2023, 2024; Zhou et al., 2024c, d; Schiebinger, 2021; Torregrosa and Garcia-Ojalvo, 2021), effectively constructing these landscapes using single-cell omics data remains the major challenge. In recent works (Zhu and Wang, 2024; Zhu et al., 2024a), the authors utilize RNA velocity to construct a vector field from the snapshot scRNA seq data and then compute the potential landscape based on the Boltzmann distribution-like relations proposed by Wang et al.. To be precise, the landscape is characterized by the expression
@@
where $p_{\text{ss}}$ represents the steady-state probability density function (PDF) that satisfies the steady-state Fokker-Planck equation: $-\nabla\cdot(p_{ss}\mathbf{b})+\frac{\sigma^{2}}{2}\Delta p_{\text{ss}}=gp_{\text{ss}}$. For the temporal scRNA-seq data, following (Zhang et al., 2025; Tong et al., 2024b), it enables a natural inference of the time-evolving potential energy landscape by leveraging the learned log-density function. Specifically, one can define the landscape at time $t$ as
$U(\mathbf{x},t)=-\frac{\sigma^{2}(t)}{2}\log p(\mathbf{x},t).$
@@
5.4 Other Challenges and Further Directions
Integrating multiomics data is critical for comprehensively characterizing cell-cell interaction dynamics and regulatory mechanisms (Stein-O’Brien et al., 2021). Furthermore, aligning both temporal and spatial scales within time-series spatial transcriptomics data (e.g., using rigid/non-rigid spatial transformations, and latent space) presents a significant challenge. Additionally, the application of SDEs in modeling spatial transcriptomics data, and subsequently constructing spatiotemporal developmental landscapes, represents an important avenue for further exploration.
6 Discussion and Conclusion
@@
Due to the limited scope of the current review, several important aspects of the dynamical models of scRNA-seq have not been discussed thoroughly here and remain for further exploration. First, lineage tracing plays a critical role in understanding cellular history and developmental trajectories and integrating such data into trajectory inference could provide deeper insights into cellular fate decisions (Wagner and Klein, 2020; Weinreb et al., 2020; Ventre et al., 2023). Second, incorporating gene regulatory networks (GRNs) (Pratapa et al., 2020; Van de Sande et al., 2020; Zhang, 2024; Stumpf, 2021; Akers and Murali, 2021) into spatiotemporal trajectory inference is an exciting avenue for future research, as it could enhance the understanding of the regulatory mechanisms driving cellular transitions. Finally, the concepts of dynamic network biomarkers (DNBs) and critical transitions (Liu et al., 2017; Zhang et al., 2024c; Chen et al., 2018b; Han et al., 2022) are promising for understanding cellular fate shifts, particularly in disease progression and cellular development.

### v2

@@ 5.3 Reconstructing Waddington Developmental Landscapes
Waddington’s landscape metaphor is a widely recognized framework to depict the cell fate decision process. This conceptual model suggests that metastable cellular states are analogous to wells within a potential landscape, and transitions between these states can be understood as movements or “hops” between these potential wells. While the development of such potential landscapes has been extensively explored Ao (2004); Zhou and Li (2016); Shi et al. (2022); Zhao et al. (2024); Li and Wang (2013); Wang et al. (2010); Li and Wang (2014); Bian et al. (2023, 2024); Zhou et al. (2024, 2024); Schiebinger (2021); Torregrosa and Garcia-Ojalvo (2021), effectively constructing these landscapes using single-cell omics data remains the major challenge. In recent works Zhu and Wang (2024); Zhu et al. (2024), the authors utilize RNA velocity to construct a vector field from the snapshot scRNA seq data and then compute the potential landscape based on the Boltzmann distribution-like relations proposed by Wang et al. Wang et al. (2008). To be precise, the landscape is characterized by the expression
@@
where $p_{\text{ss}}$ represents the steady-state probability density function (PDF) that satisfies the steady-state Fokker–Planck equation: $-\nabla\cdot(p_{ss}\boldsymbol{b})+\frac{\sigma^{2}}{2}\Delta p_{\text{ss}}=gp_{\text{ss}}$. For the temporal scRNA-seq data, following Zhang et al. (2025); Tong et al. (2024), it enables a natural inference of the time-evolving potential energy landscape by leveraging the learned log-density function. Specifically, one can define the landscape at time $t$ as
$U(\boldsymbol{x},t)=-\frac{\sigma^{2}(t)}{2}\log p(\boldsymbol{x},t).$
@@
5.4 Challenges and Further Directions
Integrating multiomics data is critical for comprehensively characterizing cell–cell interaction dynamics and regulatory mechanisms Stein-O’Brien et al. (2021); Cang and Zhao (2024); Demetci et al. (2022); Zhou et al. (2023); Cao et al. (2022); Xia et al. (2023); Gao et al. (2024); Tang et al. (2024); Lahat et al. (2015); Liu et al. (2023). Furthermore, aligning both temporal and spatial scales within time-series spatial transcriptomics data (e.g., using non-rigid or non-linear spatial transformations, and latent space) presents a significant challenge. Additionally, the application of SDEs in modeling spatial transcriptomics data, and subsequently constructing spatiotemporal developmental landscapes, represents an important avenue for further exploration.
6 Discussion and Conclusions
@@
When presenting existing approaches, we chose various types of dynamical systems modeling approaches as the main focus, providing a systematic overview of their applications across different contexts. Specifically, we examine the utility and limitations of dynamic modeling techniques in four distinct data scenarios: (1) single time-point scRNA-seq data, (2) multi time-point scRNA-seq data, (3) single time-point spatial transcriptomics data, and (4) multi time-point spatial transcriptomics data, i.e., spatiotemporal single-cell data. For each data type, we explore how different modeling paradigms—ranging from discrete Markov chain models to continuous Ordinary Differential Equations (ODEs), Stochastic Differential Equations (SDEs), and Partial Differential Equations (PDEs)—can be used to study the underlying biological processes embedded in the snapshot and high-dimensional nature of sequencing data.
For static single snapshot scRNA-seq data, where explicit temporal information is unavailable, top-down discrete-state models such as Markov chains have been widely adopted to infer latent cell state transitions and developmental trajectories. We also address how bottom-up mechanism models like RNA velocity methods provide insights into modeling complex state transition dynamics. When explicit temporal resolution is introduced as in time-series scRNA-seq datasets, continuous dynamical models become more suitable. Incorporating the dynamical optimal transport (OT) theoretical framework, the Fokker–Planck PDE-based models allow us to track the evolution of cellular states over time. The integration of spatial transcriptomics adds another dimension to the modeling challenge. In the case of multi-time-point spatial transcriptomics, we review emerging approaches that combine dynamical OT and geometric transformation to simultaneously account for trajectory inference and batch correction across time points.
Beyond summarizing the mathematical aspects of existing models, we also provide a forward-looking perspective on potential future directions. For instance, integrating both discrete and continuous models, as well as cellular interaction effects, could provide new insights to handle more realistic dynamics with increasing biological interpretability. Additionally, the use of multi-modal omics could also enhance the resolution of dynamical models. To sum up, by examining the intersection of dynamical systems theory and single-cell data modeling, this review provides conceptual and methodological insights that may inspire the development of novel algorithms to dissect the spatiotemporal dynamics underlying single-cell sequencing data.
Due to the limited scope of the current review, several important aspects of the dynamical models of scRNA-seq have not been discussed thoroughly here and remain for further exploration. Firstly, lineage tracing plays a critical role in understanding cellular history and developmental trajectories, and integrating such data into trajectory inference could provide deeper insights into cellular fate decisions Wagner and Klein (2020); Weinreb et al. (2020); Ventre et al. (2023). Secondly, incorporating gene regulatory networks (GRNs) Pratapa et al. (2020); Van de Sande et al. (2020); Zhang (2024); Stumpf (2021); Akers and Murali (2021); Zhao et al. (2024); Yang (2025) into spatiotemporal trajectory inference is an exciting avenue for future research, as it could enhance the understanding of the regulatory mechanisms driving cellular transitions. Lastly, the concepts of dynamic network biomarkers (DNBs) and critical transitions Liu et al. (2017); Zhang et al. (2024); Chen et al. (2018); Han et al. (2022) are promising for understanding cellular fate shifts, particularly in disease progression and cellular development.
@@
[…]
Conceptualization, Z.Z., P.Z. and T.L.; investigation, Z.Z., Y.S., Q.P., P.Z. and T.L.; writing—original draft preparation, Z.Z., Y.S., Q.P. and P.Z.; writing—review and editing, Z.Z., Y.S., Q.P., P.Z. and T.L.; visualization, Z.Z., Y.S., Q.P. and P.Z.; supervision, P.Z. and T.L.; funding acquisition, P.Z. and T.L. All authors have read and agreed to the published version of the manuscript.
[…]
This work was supported by the National Key R&D Program of China (No. 2021YFA1003301 to T.L.) and National Natural Science Foundation of China (NSFC No. 12288101 to T.L. & P.Z., and 8206100646, T2321001 to P.Z.).
[…]
The authors declare no conflicts of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
scRNA-seq single-cell RNA sequencing
SDE Stochastic Differential Equation
ODE Ordinary Differential Equation
PDE Partial Differential Equation
OT Optimal Transport
ICA Independent Component Analysis
PCA Principle Component Analysis
MST Minimum Spanning Tree
MCE Markov Chain Entropy
GPCCA Generalized Perron Cluster Cluster Analysis
EM Expectation Maximum
VAE Variational Autoencoder
RKHS Reproducing Kernel Hilbert Space
CNF Continuous Normalizing Flow
FM Flow Matching
CFM Conditional Flow Matching
SB Schrödinger Bridge
RUOT Regularized Unbalanced Optimal Transport
GWOT Gromov–Wasserstein optimal transport
FGWOT Fused Gromov–Wasserstein optimal transport
{adjustwidth}
[…]